TCCON > Sites > East Trout Lake > Instrument History > Site Visit 2017-10-05 - 2017-10-12

Site Visit 2017-10-05 - 2017-10-12

Shopping list - Rona Prince Albert - Purchased Oct 3 for pickup by Joseph Mendonca

 

Shipping List (ship everything Sept 28?)

  • IFS alignment kit. (shipped by Purolator ground, arrive Oct 3 - 331178017116, 331178017108, 331178017090, 331178017082)
  • Heat pads for suntracker.(tracking  331179275374, ETA Oct 04.)
  • Paroscientific. (Oct 3 - tracking 331178017116, 331178017108, 331178017090, 331178017082)
  • Disk.(Oct 3 - tracking 331178017116, 331178017108, 331178017090, 331178017082)
  • Vent (for pump room). - in alignment kit - (Oct 3 - tracking 331178017116, 331178017108, 331178017090, 331178017082)
  • IFS tool kit.  (Oct 3 - tracking 331178017116, 331178017108, 331178017090, 331178017082)

 

To do List

1. Check pump line to IFS for leaks.

  • Disconnect every junction between the pump and the IFS, check to ensure that the o-rings are clean and/or if they need a *tiny* bit of vacuum grease.
    • To apply vacuum grease, put a small amount on the o-ring, and then wipe off with a kim wipe.
  • Tighten connections and test pump down.
  • Pump maintenance? Hopefully unnecessary.

2. Align IFS

3. Align sunlight into aperture wheel (input field stop)

  • Make sure that solar disk is centered on the apeture (to do on sunny day), near noon. ***Make sure that you balance the lid of the source compartment on its side to maintain the weight of the lid while doing this***
    • https://tccon-wiki.caltech.edu/Caltech-built_Containers/Hardware/Aligning_the_Bruker_125HR#IR_Alignment

4. Calibrate Zeno pressure sensor.

  • Bring Paroscientific Digiquartz pressure standard to do calibration.
  • Need to put IFS in hold inorder to keep taking pressure measurements with Zeno pressure sensor.

5. Install vent(s) in pump room.

  • I'm not sure where the extractor fan is located but you'll want to install the inlet vent in the opposite corner. Ideally you will have the extractor fan located up high in one corner removing the heat and then the inlet should be placed down low on the opposite corneer. This will be complicated by the equipment (fridge) and you may need to use some duct work to avoid equipment.See the link for a rought schematic. Sites/East Trout Lake/Instrument History/Site Visit 2017-10-05 - 2017-10-12/IMG_20171003_132152164.jpg 
  • Here's a rough diagram. Note that Vent 2 may have to move up or somehow out of the way of the existing equipment. Vent 1 should be close (~1 foot up or so) from the floor. Please don't forget to check for studs and electrical cables before you drill.

pump_room_vent_proposal.jpg

  • Once you've descided on the location of the inlet vent use a stud finder to locate where the studs (wooden 2x4" vertical beams) are located. Typically they're spaced 16" apart although sometimes 24". Locate two studs and mark your vent to be installed between them
    • Make sure you're not near any AC. The stud finder works by taking a reference measurement when you hold the button, then any values measured above this correllate to a stud or AC. Hence if you calibrate the sensor over a stud or AC you won't find any studs/AC. Be careful you calibrate it where there's no stud or AC. Check for outlets nearby that could provide information about where the electrical wiring could be.
  • Drill a pilot hole at the location marked for the vent. Drill from the inside of the building to the outside. You want to make it all the way through so you can drill the 4" hole from both sides. 
  • Place the 4" hole saw bit on the mandrel (center drill bit). Be careful not to cross thread it and do not thread it all the way on. Stop threading it on about a 1/4 turn before it locks and lock in the pins. 
    • you want the pins to hold the hole saw, not the threads. 
    • when you drill the hole it will most probably go through like butter as you're not going through anything solid. 
    • In general you want to keep pressure on the hole saw bit as you drill. If you don't it might chatter around and not stay centered. 
  • Drill the first 4" hole from outside of the lab. It's best to drill half way through the wall and then go back inside the lab and drill the remaining half. The wall might be so thin that you actually go through in one pass from the outside. 
  • Now you can install the vents
  • Remove flaps from vent that will sit outside (the curved one to prevent rain from getting in).
  • Put mesh on vent to prevent bugs from getting into lab.
  • If necessary place flexable 4" ducting to route the inlet vent into a convenient place (opposite corner of exhaust fan), see Sites/East Trout Lake/Instrument History/Site Visit 2017-10-05 - 2017-10-12/IMG_20171003_132152164.jpg 
  • Test with pump room door closed to ensure that the extractor fan can indeed extract air. Place hand on inlet and outlet, inside and outside lab.
    • compare flow to that with pump room door open and to pump room door and external door open. If not satisfactory perhaps we install second inlet. It is likely that we'll need a second in

6. Wrap heaters, themorstats and insulation around pedestal.

  • Glue heaters and thermostats to pedestal using RTV adhesive.
  • Wrap reflective insulation around heaters and pedestal.
  • Wrap Tyvek outside the insulation and fasten with tuck tape.
  • Run wires through hole with the zeno wires?

7. Re-glue ST heat temp sensor on elevation plate.

  • The ST heat temp sensor in the suntracker enclosure has come off and is begin held down by something.
  • RTV adhesive to do this.

8. General lab clean up.

  • Pack up and ship back tracker control boxes, enclosure control box, and norhof tube.
  • Organize and put tools away.
  • Ensure nothing is left on the IFS.
  • Generate inventory of lab items.
  • Take detailed photos of entire lab and pump room and outside of building 360 degrees (walk around building), 60 photos on Sony a5000. 

9. Check InSb detector fill tubing.

  • I suspect that the tubing from the Norhof is not properly seated on the top of the InSb detector (or something has come loose).
  • Check the tubing and perhaps fasten the black "puck" back down on the InSb detector.
  • Pull puck out of the InSb to check whether the tube that distributes the LN2 inside the detector is functional/attached/cracked/correctly oriented, etc. Also check to ensure that the temperature sensor is not in the direct flow of LN2 (away from the hole in the tube).

10. Make foam box for O2 sensor or possibly repostion.

  • Need foam for this. Might be some left over at ETL from IFS install (but might have been discarded). Could ship or purchase some.
  • Least important.
  • Could move to main lab room.

11. Fix HERC monitor.

  • Check cables to HERC monitor.
  • Might need to replace with new monitor. <-- New monitor was shipped last week. It's supposed to arrive on Monday, Oct 2, but Greg may not be able to pick it up before you leave.

12. LN2 generator maintenance <-- Greg to handle. 

13. InSb jacket check.

  • The InSb isn't staying cold for very long. Fill the detector and see if there's condensation, or if the outside of the jacket is very cold. If so, we'll need to bring the InSb detector back to Toronto to pump out the jacket.
  • Please show Greg how to reinstall the InSb detector. Do *not* have him adjust any of the mirrors. Ensure that the detector is peaked before removing it.

Safety tips

  • Before connecting to the IFS or tracker with a laptop, shut down IFSloop:
    • sw status shutdown quickly
  • To start IFSloop back up after a shutdown:
    • IFSloop start
  • To start IFSloop after merely exiting:
    • IFSloop
  • Remember you'll have to reconnect the herc to the suntracker upon restart after controlling the tracker with a laptop. This can be done from IFSloop, I believe. This is the sequence the Norton has used in the past. You may want Norton or Debra to run the commands if you're not comfortable doing so yourself.
    • 'SW Status Reinit' followed by a pause, and then 'SW Status Time Warp' to open the enclosure.
    • Sun Tracker Connect - to get it out of stale (if the stale counter is rising)
    • Sun Tracker Init - to retry the init (you can watch the tracker rotate its mirrors)
    • Sun Tracker Track to Programmed Sun Position
    • Sun Tracker Track by Diode
  • If you want to use the switches on the suntracker enclosure to manually open the suntracker enclosure do the following
    • Stop HERC control of cover
    • Disconnect measurement computing (blue boxes on wall above Herc) power supply
    • ensure that the herc control cable is disconnected from the suntracker enclosure control box.
    • Now you can use the switches on the suntracker enclosure to open/close the cover.
  • Before reconnecting the Herc to the suntracker control box.
    • Leave the switch on the suntracker enclosure in the middle "auto" position
    • Connect the Herc control cable to the suntracker enclosure control box
    • Power on the Measurement Computing blue relay/A-D modules (located above the Herc on the wall).
  •  

 

October 5, 2017

  • Today Jacob, Sebastien, and Joseph arrived at the site and unloaded all of the equipment. By the time they were done it was dark outside so they decided to go back to the resort. 
  • Greg started the LN2 genrator today. Joseph did the checks before leaving the lab. The dewar was about 3/4 full. 

Back to top

October 6, 2017

  • Jacob, Sebastien , and Joseph arrived at the lab around 7:30 am local time.
  • HERC was filling LN2 when arrived at the lab. ***Important*** Joseph recorded video of liquid N2 coming out of top while LN2 was filling (will be attached eventually).
  • The HERC screen is not working. Disconected from the HERC and connected to Sebastien computer. Confirmed that the screen was not working properly. Not a problem with the HERC.
  • Joseph was able to ssh into the HERC.
    • But typed "ph" which caused the screen to freeze. Joseph closed the screen and tried to reconnect to the HERC but no reponse. Orfeo tried to ssh into the HERC and was not succesful.
    • Debra gave Joseph instructions on how to power cycle the HERC.
    1. Turn off switch on silver box.
    2. Disconnect power cord from HERC.
    3. Wait a few seconds.
    4. Reconnect power and turn switch on.
    • HERC was on again and lights were green.
    • Joseph was able to SSH into the HERC. Debra ssh into the HERC to determine status.
    • HERC stayed in IFSloop all day.
  • Heaters for the suntracker pedestal.
    • First glued the heaters and themosats to pedestal using RTV adhesive.
    • Then wrapped all the insulation around the heaters with the tuck tape.
    • Wrapped the instulation with the tyvek using the tuck tape.
    • Ran the heater extension cords from the roof through the hole where the Zeno wires come in.
  • Joseph re-glued the ST heat temp sensor on the elevation plate using the RTV adhesive.
  • Sebastien started to put together the telescope for the alignment.
  • Sebastien figured out how to control the camera that will be used for the alignment with Joseph cell phone because the HERC screen is not working.
  • Cleaning the pump line.
    • When cleaning the pump line they found that there was a lot of dark grey dust lineing coating the inside of the vacuum tube. The closer to the pump the more dust.
    •  
    • They cleaned the tubing with kim wipes. The o-rings and seals were cleaned with km wipes and methanol. Vacuum grease was put on the orings after they were cleaned.
    • Joseph was able to run power Pump on to turn on the pump.
      • In a few seconds the pump pressure was at 0.5 as shown on the HERC. After about a minute it was at 0.45.
      • ***Important*** both Jacob and Joseph herd a grinding noise when the pump was turned on. Orfeo thinks that the scrolls a grinding against something causing the grey dust. 
      • Joseph turned the pump off using power Pump off.  The pump pressure went back up.
  • Jacob ran the paroscientific (placed it on the roof near the suntracker) for about 2hrs today while IFSloop was running.
  • Jacob numbered the foam covering the IFS and took pictures so that it will be easy to put foam back on the same spot after it is removed for the alignment.
  • Joseph filled the chiller with deionized water. The chiller was about 1/5 empty.
  • Jacob vacumed the lab.
  • Joseph did the checks on the LN2 generator the dewar is almost full.
    • The water filter is in the yellow range now. ***Important*** Greg will have to change it in the near future. He will require instructions for that.

 

Tomorrow

  • Test the heat pads on pedestal by pluging in to Belkin in the morning (hopefully it will be cold enough).
  • Tomorrow main task is the IFS alignment.

Back to top

October 7, 2017

Before_alignment.jpg

  • Got to lab at about 7:30 am local time.
  • HERC was filling LN2 and pump was on Pump_p = 0.85.
  • Jacob and Joseph plugged in one of the pedestal heaters. It was on, belkin showing about 940 watts. Left on for 5min until Jacob thought he smelled something. Joseph didn't smell anything but unpluged the heaters just incase.
  • Took the foam off the IFS and vacuumed it.
  • HERC began pumping down IFS. IFS_p is staying at 0.81 hpa while pump_p is staying around 0.75 hpa.
  • FYI: I've attached the daily-etl-checklist.pdf file we discussed today. You might want to print out a copy (at least p4/Table 1) and walk Greg through it.
  • Waited until instrument was evacuated, we let it take a few scans and issued the sw status shutdown quickly.
    • Suntracker was still open and tracking. Used Bruker softwear to stop it.
    • Had to close the cover manually. Used the switch at the bottom of the tracker. Switched from center postion to close (move switch to the right. ***Important*** need to put switch back to center postion so the HERC can control the suntracker
  • Alignment 
    • We first focused the telescope.
    • The telescope with the camera was to big to fit int the scanning arm compartment. Had to take off the white cylinder infront of the telescope. It was able to fit in.
    • To test the camera telescope setup we looked at how the apetures were aligned. The image of the apetures was transmitted to Joseph phone and we coulds see the apetures.
    • We took the telescope out of the scanner compartment and placed it on the tripod that was next to interferometer compartment. 
    • We were able to mount the external laser on the source compartment. The pareoscope was in the interformeter compartment.
    • Placed 3 pices of optical paper on and a diffuser (1500) in front of the feild stop. The HCl cell was removed but the holder left in the same place. The specel reducer was used used to vibrate the optical paper only.
    • Fringe pattern was visible on the phone. The phone was used to take pictures of it. At 25cm OPD we could crealy see that the center circle was not centered.
    • We took pictures of the fringe pattern at OPD = -5,0,5,25,40,60,70 cm.  
    • Calculated the laser A and B ME. Before doing any alignment.
    •   ME @ ZPD ME @ max OPD
      LasA 0.72 0.61
      LasB 0.70 0.55
    • The flat mirror was adjujsted so that the fringe pattern would be centered at OPD = 25 and 60 cm. Going back to OPD = 5,0-5 and looking at the fringe pattern, the adjustments did not help. Ths means shaer misalignment.
    • The fixed cube corner was unlocked so that it could be adjusted. This meant that the plate between the source and interferometer compartment had to be removed.
    • The fixed cube corner was adjusted so that we could see the circle at OPD = -5,0,5 cm. Then we went to OPD = 25 and 60 cm and adjusted the flat mirror (only slightly) to center the circle. We went back to OPD =-5,0,5 to see if things changed. The fringe pattern was still centered. We took pictures of the fringe pattern at OPD = -5,0,5,25,40,60,90 cm.
    • Next we aligned the feild stop with the internal apeture. Going through different size apeture settinngs to make sure they are aligned.
    • Finally we did the laser alignment. The ME stayed the almost the same from OPD = 0 to 90 cm.
    •   ME @ ZPD ME @ max OPD
      LasA 0.66 0.60
      LasB 0.62 0.57

October 8, 2017

  • Alignment
    • Centered the lamp beam on the feild stop. Confirmed it was focused on the exit apeture.
    • Peaked both Ingaas and Insb detectors. 
    • Took multiple HCl spectra. Tried to analyse in LINEFIT14 but gives and error "significant spectral detuening required". At first we thought this was a problem with the Laser alignment. So we aligned it multiple times but ME stayed the same. Debra thinks this is because the cell test was taken while the instrument was vented. Sent input files to Frank Hase to try and get it to work.
  • Put suntracker in auto. The dome opened again.
    • Started IFSloop using "IFSloop start".
    • Put in hold using "sw status hold". This closed the dome. But after about an hour gave error message that can't comunicate with the dome.
    • After trying LN2 filling did a "sw status shutdown quickly" to get out of IFSloop, so that we can open the dome manually tomorrow and use Bruker webpage control the suntracker.
  • Trouble shooting LN2 filler.
    • Took out the LN2 "hockey puck" from Insb. Inpected the hose and the temp sensor, both look normal.
    • Put the puck back on the Insb and made sure the velcro was tight.
    • Issued "sw status Insb Fill" to fill the Insb. LN2 was pumped into the insb and temperature dropped slowly. Intially a small amount of liquid N2 could be seen coming out but stopped after a couple minutes. HERC temp was at -186 after about 20min, but got error message that the Insb was not filled. Will try again.
  • Jacob was able to find the studs and electrical wiring on the wall where vent two is suppose to go.
    • He marked out a space above and to the left of one of the aresol pumps. Debra is concerned that if we put the vent there cold air will be on the aresol pump line. Wiating for envirnoment Canada to aprove of vent location.
    • Got the drill from Greg. Joacb was able to find all the drill parts to make the hole in the container.
  • Jacob worked on building a foam box for the O2 sensor.

Back to top

October 9, 2017

  • Cold day this morning so we pluged the heaters back in. We are monitoring the wattage using the Belkin. Opened the suntracker dome to be able to touch the pedestal.
    • When we plugged in the heaters (one set) it came on.
    • Inside of the pedestal felt a little bit warmer after about 10 min. The heaters also shut off.
  • Jacob finished building the foam box and put it around the O2 sensor.
  • IFS Alignment 
    • Frank Hase got back to us. His advice was to put the "significant spectral detuning value" (-2.9e-6) in the LINEFIT input file for the parameters "take species into account(T/F),column scaling factor, spectral scaling factor of species - 1" change all line from ".true.,1.0,0.0" to .true.,1.0,-2.9e-6. ME at maxOPD was 1.1.
    • Did the alignment again adjusting fix cube corner. Coaligned the apetures. Aligned the laser. Ran a  50 scan vented HCl cell test. From Sebastien: "When we did the alignment again I undid the change in the screw of the horizontal axis of the fixed corner cube to see if I had been compensating an angular misalignment with it. But we confirmed that we were seeing a shift of the fringe pattern on either side of ZPD that could not be compensated by adjustments of the flat mirror. We did the same check with the vertical position."
    • Using Linefit 14.5, the ILS looks good. The ME at max OPD was about 0.986. Did different reg factors (for TCCON extended ILS) and found the ME stayed the same (around 0.986). 
    • Laser ME results:
    •   ME @ ZPD ME @ max OPD
      LasA 0.63 0.52
      LasB 0.66 0.52
  • Align Sunlight into apetrure wheel
    • It was sunny around noon while doing the IFS alignment.
    • Used Bruker softwear to control tracker. Had to configure fire wall and Java as done in last site visit (https://tccon-wiki.caltech.edu/Sites/East_Trout_Lake/Instrument_History/Site_Visit_2016-09-19_-_2016-09-30). Was able to track sun at 1:30 pm local time.
    • Blocked sunlight from going into detector.
    • Was able to see that solar beam was not centered on 4mm apeture. So centered it (using flat mirror) while lid was placed vertically in the middle on top of source compartment.  
  • Installing vent
    • Jacob drilled a guide hole through the wall, as well as the hole from the inside.
    • The metal sheet on the outside of the container was very hard to drill through. Gregs drill broke whil trying to drill wide hole from outside. Jacob tried to get it off with wire sheers and a chisel
    • After lunch we borrowed Bob's (owner of pine grove resort) drill. Was able to drill through, but it took a VERY long time.
    • Eaiser tor drill through the the wood but that took some time as well.
    • Installed the vent. Used the angle gringer to cut metal sheet to size of the hole.

Back to top

October 10, 2017

alignment-2.jpg

  •  IFS Alignment
    • Used IFSloop to evacuate the instrument to do an evacuated HCl test. Typed "as Status Pumpdown". This took about 2hrs.
    • LINEFIT14.5 results were poor. ME at max OPD (45 cm) was 0.92. Posibly something moved during  evacuation of the instrument.
    • Did anonther cell test vented and ME was at 0.96 at OPD=45cm. Sebastien thinks that maybe cube corner spring screws could be a bit tighter.
    • Re-aligined again. This time after fixing the shear (iterating back and forward betwen +5, near zpd and -5 cm), went back to 25, 45, and 90 cm and looked at fringies with apetures 4, 2.5, and 1 mm adjusting flat mirror to center fringes as much as possible. After tightening the spring screws that keep the fixed corner mirror in place we looked at the fringes at  -5, near zpd, +5, 25, 45, and 90 cm with apetures of 4, 2.5, and 1 mm to make sure fringes still centered, which they were. co-aligned the apetures.
    • Did an HCl cell test (50 scans). LINEFIT 14.5 result was 0.992 ME at OPD=45cm.
    • Did the laser alignemnet and detector peaking. When detector peaking original postion gave bset results.
    • Laser ME:
    •   ME @ ZPD ME @ max OPD
      LasA 0.72 0.67
      LasB 0.62 0.62
    • Another HCl test results (after laser peaking) was 0.992 ME at OPD=45cm, but the column scale factor was closer to 1 vs previous test.
    • Evacuated the Bruker using the following commands in IFSloop (after disconnecting from OPUS and the Bruker webpage)
      • IFSloop start (to satrt IFSloop)
      • sw status hold (to put container in hold)
      • IFS Direct VAC=0 (to put Bruker in standby, extra check even though we put in standby using the Bruker webpage before closing it)
      • power pump on (to turn on the the pump). Took a few seconds for the pump to get down to ~0.3hpa. Waited about 15 min before issueing next command.
      • IFS Direct VAC=1 (to evacuate the Bruker). This took about 20min for the Bruker to evacuate. While the Bruker is evacuating it goes into error. Once it get below 1 hpa it will beep and the error light will go away. After about 20min the pump was at ~0.5hpa.
      • IFS Direct VAC=0 (to put the Bruker back into standby). This will close the flaps which will make a loud clunk noise.
      • power pump off (to shut pump off). Waited until pump got back down o ~0.29 hpa to shut pump off.
      • sw staus shutsown quickly (to shut down IFSloop before using the Bruker webpage again).
    • Evacuated HCl cell test analysis showed ME of ~0.94 at OPD=45cm.
    • Vented instrument and did another HCl cell test. Got ME of 0.992 ME at OPD=45cm. Don't think anything moved this time unless it moved back when venting the Bruker.
    • Noticied that OPUS parameter "atmospheric compensation" was selected for all test. Redid the vented and evacuated test again eith this removed but LINEFIT 14.5 results were the same.
    • Tried to alignment againg. This time tried to do alignment without changing the vertical screw much but it was not working so had to change it. Went through the previous steps. Fringes looked centered. Hard to co-aligin the apetures this time, lighting problems, could not get them lit as well as the previous time. 
      • Did a evacuated cell test results were not good.
      • Image of the source was not centered on the exit apeture well. Used the flat missor infron of the exit apeture to center image and did another HCl test. ME wnet up by 7%.
  • Align Sunlight into apetrure wheel
    • Sunny this morning so used Bruker to control suntracker to check that sun was centered on the entrance apeture, which it was not. It was off by about 1/3 of the diameter of the beam.
    •  At about 12:15 noon used aligned the beam again.
    • At about  2:30 pm (local time) checked again that the beam was centered on the apeture and it was.
  • Installing vent
    • Jacob was able to mark out where to install the second vent.
    • He drilled a guide hole through the wall, as well as the hole from the inside.
    • It took a lot of time and work it get the metal sheet off the outside of the building.
    • Not enough time to finish up so he plugged the hole. 
      • Just have to drill the hole from the outside and install the vent.
  • Packing
    • Packed up suntracker controller, 2 dome enclosure boxes and norhof fill hose in one box.
    • Packed up the IFS tool kit.
  • LN2 generator
    • Generator for running for 5 days. Black indicator is at full and two orange indicators are at full and 3/4.
    • Tried filling Norhof with LN2 but nothing was coming out. Black indicator would just fall to zero when trying to fill. When stopped filling it would got abck to full.
    • Orfeo said that the pressure on the black part of the dewar should be between 7-10 psi. It was at 0.
      • Orefeo sugested turining the nobe next to pressure indicator to somewhere between 7-10 psi. Which we did.
      • Later in the evening I checked that hoses connected to the cold head. One is suppose to be hot and the other cold. Both were room temp.
    • Greg arrived late in the evening. he shut off the LN2 generator. 

Back to top

October 11, 2017

  • IFS Alignment
    • Discussion with Debra about results obtained by LINEFIT 14.5 led to us not changing anything about the alignment this morning.
  • Packing
    • Packed up the alignment kit.
  • IFSloop
    • started IFSloop.
    • sw Status Renint. To start taking measurments again. All stale counters were zero. Took about 3min to communicate with Bruker again.
    • Tried issuing commands to suntracker but not responding. Turned tracker box on and off. Did a suntracker reinit and was able to communicate with tracker.
    • Measurements resumed.
  • Insb Fill
    • Tank depth is low but still enough to fill Insb.
    • When filling liquid LN2 was coming out off top.
    • Waited until norhof hose thawed. Put the red fill part into the clear part better and put puck back on top of insb. Extra velcro to keep puck fastened down better.
    • Didn't get a chance to do insb fill before leaveing.

Back to top 

Shopping list suggestions (next visit)

  • Outside light bulb/fixture
  • Extension cord, 50 ft?
  • Wall clock (one in there did not seem to be working anymore)
  • 9 V batteries, for stud finder/multimeter
  • M purple gloves
  • Ziploc bags (sandwich, gallon?) for storing small parts

 

Longer term suggestions

  • IFS toolbox to leave on site. It seems expensive to ship the entire (heavy) box back/forth. We only used a few of the tools anyways.
  • Dedicated laptop (semi-rugged?) with OPUS, linefit, etc. installed. Could possibly be shared with another project.

 

Back to top

External links

Photos taken with my phone (Jacob) - https://photos.app.goo.gl/0ibGTTQ3xpfhh2O42 

 

Back to top

 

Tag page

Files 17

FileSizeDateAttached by 
 20171006_143019.jpg
No description
3.89 MB16:47, 7 Oct 2017dwunchActions
 20171006_143846.jpg
No description
3.86 MB16:47, 7 Oct 2017dwunchActions
 align.pdf
Frank Hase Karlsruhe 125HR Alignment Instructions
57 kB11:15, 3 Oct 2017orfeoActions
 bruker_125_align.pdf
John Robinson NIWA 125HR Alignment Instructinos
674.17 kB11:15, 3 Oct 2017orfeoActions
 Cleaned pump line.jpg
No description
3.86 MB10:35, 13 Oct 2017mendoncaActions
 Cross Hairs Fringe alignment.pdf
Fringe Alignment Guide
33.23 kB06:35, 4 Oct 2017orfeoActions
 Cross Hairs Fringe alignment super wide spacing.pdf
Fringe Alignment Guide Super Wide
20.95 kB06:35, 4 Oct 2017orfeoActions
 Cross Hairs Fringe alignment wide spacing.pdf
Fringe Alignment Guide Wide
23.75 kB06:35, 4 Oct 2017orfeoActions
 daily-etl-checklist.pdf
ETL Daily Checklist - https://www.overleaf.com/read/vhnyqpdnjhqw
292.52 kB17:05, 7 Oct 2017dwunchActions
 ETL_Oct2017_pressure_comp.pdf
Pressure sensors comparison
497.69 kB07:37, 20 Oct 2017jhedeliuActions
 Eureka_alignment.pdf
Sebastien Roche Alignment Instructions
4.06 MB07:21, 4 Oct 2017orfeoActions
 Grey ash in pump line.jpg
No description
3.89 MB10:35, 13 Oct 2017mendoncaActions
 IFS125_alignment_Wollongong_20150709.pdf
David Griffith Wollongong 125HR Alignment Instructions
2.07 MB11:15, 3 Oct 2017orfeoActions
 IMG_20171006_144230.jpg
No description
4.08 MB16:47, 7 Oct 2017dwunchActions
 IMG_20171006_145816.jpg
No description
3.88 MB16:47, 7 Oct 2017dwunchActions
 Procedure_for_aligning_IFS125.doc Preview
Debra Wunch Caltech 125HR Alignment Instructions
46.5 kB11:17, 3 Oct 2017orfeoActions
 pump_room_vent_proposal.pdf
No description
16.03 kB16:54, 7 Oct 2017dwunchActions